# MiLe Loss: a New Loss for Mitigating the Bias of Learning Difficulties in Generative Language Models

### v6

@@ 4.2 Experimental setup
BoolQ
HellaSwag
LAMBADA
OpenBookQA
PIQA
SIQA
StoryCloze
Winogrande
Avg
@@
0-shot Cross-Entropy Loss
57.52
40.73
39.10
30.60
67.08
40.79
63.55
53.75
49.14
Focal Loss
58.35
41.17
40.09
32.80
67.25
41.91
63.07
51.70
49.54
MiLe Loss 59.57 41.27 41.34
30.00
67.25
41.61
63.60 54.78 49.93
1-shot Cross-Entropy Loss
54.22
40.86
37.16
30.40
67.85
41.66
62.69
53.04
48.48
Focal Loss
53.64
41.04
37.88
32.20
67.14
44.27
62.16
52.64
48.87
MiLe Loss 55.23
40.90
38.75
32.00
67.68
43.35
63.23 55.88 49.63
5-shot Cross-Entropy Loss
50.89
41.06
36.27
28.80
67.68
43.39
62.37
50.99
47.68
Focal Loss
48.10
41.80
38.50
31.40
67.19
46.01 63.01
52.09
48.51
MiLe Loss 52.29
41.53
39.05
28.80
67.41
45.39
62.85
54.06 48.92
@@
0-shot Cross-Entropy Loss
55.96
47.48
45.76
32.20
69.64
42.43
65.47
54.54
51.69
Focal Loss 62.02
47.61
46.87
33.00
69.59
42.02
65.63
55.01
52.72
MiLe Loss
56.94
47.64 47.37 33.80 70.13
41.91
66.06 55.96
52.48
1-shot Cross-Entropy Loss
54.71
47.37
42.13
34.40
69.42
44.78
65.26
56.27
51.79
Focal Loss 62.35 47.41
43.88
32.60
69.15
45.04
65.42
54.85
52.59
MiLe Loss
54.95
47.39
45.08 34.00 70.13 45.04 65.58
54.85
52.13
5-shot Cross-Entropy Loss
55.72
47.74
41.55
33.00
69.86
45.04
66.11
55.64
51.83
Focal Loss 62.17 48.00
42.87
32.00
69.75
45.60
66.01
56.20
52.82
MiLe Loss
55.38
47.78
45.00 34.00 70.13 46.26 66.22 56.83
52.70
@@
0-shot Cross-Entropy Loss 62.14
58.91
55.54
34.40
73.61
44.06
70.66
61.40
57.59
Focal Loss
59.72
59.59
55.64
36.60
73.94
43.04
70.12
61.88
57.57
MiLe Loss
60.89
59.63 57.73
35.20
73.99 44.06 71.25
61.01
57.97
1-shot Cross-Entropy Loss
59.24
58.68
53.48
37.00
73.99
47.90
70.60
60.69
57.70
Focal Loss
58.53
59.23
52.59
35.60
74.27
48.06
69.96
59.91
57.27
MiLe Loss 60.46 59.56 55.35 38.00
73.29
48.57 70.87 61.01 58.39
5-shot Cross-Entropy Loss
61.28
59.44
54.01
37.00
74.16
49.03
71.30
63.06
58.66
Focal Loss
57.98
60.10
55.91
36.80
74.05
50.0
70.44
62.90
58.52
MiLe Loss 62.20
60.06
58.16 37.80
73.61
50.67 71.67 63.30 59.68
@@ 5.2 Perplexity on the Pile Validation Set
BoolQ
HellaSwag
LAMBADA
OpenBookQA
PIQA
SIQA
StoryCloze
Winogrande
Avg
0-shot Cross-Entropy Loss
66.73
63.48
60.95
36.80
75.52
44.58
72.26
61.88
60.27
MiLe Loss 68.62 64.17 61.52 39.00
75.41
44.63 72.90 63.61 61.23
1-shot Cross-Entropy Loss
64.13
63.33
58.92
40.40
75.46
48.72
72.90
63.85
60.96
MiLe Loss 65.26 63.93 60.57
38.80
75.46 49.64
72.58
63.93 61.27
5-shot Cross-Entropy Loss
64.22
63.92
60.90
39.60
75.84
51.18
73.60
64.72
61.75
MiLe Loss 66.85 64.58 64.33 41.00
75.14
52.66 74.02 66.06 63.08
@@ 5.3 Performance of Training with More Tokens
MiLe Loss assesses the learning difficulty of each token through information entropy. Intuitively, the more tokens used in model training, the more powerful the Language Model (LM) becomes, and the output word distribution becomes more reasonable. Consequently, the assessment of the learning difficulties of tokens becomes more accurate, and thus the MiLe Loss can probably better lead the LM to tackle the bias. To validate that, with limited computational resources, we continue to pre-train the 6.7B model from 100B tokens to 200B tokens, with both Cross-Entropy Loss and MiLe Loss. Their corresponding evaluation results on all benchmarks are reported in the following Table 8. We can see that, when the number of training tokens for the 6.7B models increases to 200B, the models trained with MiLe Loss yield consistent and substantial performance improvements over those trained with Cross-Entropy Loss. Moreover, compared to training with 100B tokens, training with more tokens even helps MiLe Loss to yield LARGER performance improvements. For instance, in the 5-shot setting, with 100B training tokens, the performance improvements gained by MiLe Loss over Cross-Entropy Loss on the 6.7B models is 1.02%. Then by continuing pre-training with more training tokens, the gained improvements increase to 1.33%. The experimental results above demonstrate well that using more tokens and longer pre-training increases the benefits of MiLe Loss.

### v7

@@ 4.2 Experimental setup
BoolQ HellaSwag LAMBADA OpenBookQA PIQA SIQA StoryCloze Winogrande Avg
@@
0-shot Cross-Entropy Loss 57.52 40.73 39.10 30.60 67.08 40.79 63.55 53.75 49.14
Focal Loss 58.35 41.17 40.09 32.80 67.25 41.91 63.07 51.70 49.54
MiLe Loss 59.57 41.27 41.34 30.00 67.25 41.61 63.60 54.78 49.93
1-shot Cross-Entropy Loss 54.22 40.86 37.16 30.40 67.85 41.66 62.69 53.04 48.48
Focal Loss 53.64 41.04 37.88 32.20 67.14 44.27 62.16 52.64 48.87
MiLe Loss 55.23 40.90 38.75 32.00 67.68 43.35 63.23 55.88 49.63
5-shot Cross-Entropy Loss 50.89 41.06 36.27 28.80 67.68 43.39 62.37 50.99 47.68
Focal Loss 48.10 41.80 38.50 31.40 67.19 46.01 63.01 52.09 48.51
MiLe Loss 52.29 41.53 39.05 28.80 67.41 45.39 62.85 54.06 48.92
@@
0-shot Cross-Entropy Loss 55.96 47.48 45.76 32.20 69.64 42.43 65.47 54.54 51.69
Focal Loss 62.02 47.61 46.87 33.00 69.59 42.02 65.63 55.01 52.72
MiLe Loss 56.94 47.64 47.37 33.80 70.13 41.91 66.06 55.96 52.48
1-shot Cross-Entropy Loss 54.71 47.37 42.13 34.40 69.42 44.78 65.26 56.27 51.79
Focal Loss 62.35 47.41 43.88 32.60 69.15 45.04 65.42 54.85 52.59
MiLe Loss 54.95 47.39 45.08 34.00 70.13 45.04 65.58 54.85 52.13
5-shot Cross-Entropy Loss 55.72 47.74 41.55 33.00 69.86 45.04 66.11 55.64 51.83
Focal Loss 62.17 48.00 42.87 32.00 69.75 45.60 66.01 56.20 52.82
MiLe Loss 55.38 47.78 45.00 34.00 70.13 46.26 66.22 56.83 52.70
@@
0-shot Cross-Entropy Loss 62.14 58.91 55.54 34.40 73.61 44.06 70.66 61.40 57.59
Focal Loss 59.72 59.59 55.64 36.60 73.94 43.04 70.12 61.88 57.57
MiLe Loss 60.89 59.63 57.73 35.20 73.99 44.06 71.25 61.01 57.97
1-shot Cross-Entropy Loss 59.24 58.68 53.48 37.00 73.99 47.90 70.60 60.69 57.70
Focal Loss 58.53 59.23 52.59 35.60 74.27 48.06 69.96 59.91 57.27
MiLe Loss 60.46 59.56 55.35 38.00 73.29 48.57 70.87 61.01 58.39
5-shot Cross-Entropy Loss 61.28 59.44 54.01 37.00 74.16 49.03 71.30 63.06 58.66
Focal Loss 57.98 60.10 55.91 36.80 74.05 50.0 70.44 62.90 58.52
MiLe Loss 62.20 60.06 58.16 37.80 73.61 50.67 71.67 63.30 59.68
@@ 5.2 Perplexity on the Pile Validation Set
BoolQ HellaSwag LAMBADA OpenBookQA PIQA SIQA StoryCloze Winogrande Avg
0-shot Cross-Entropy Loss 66.73 63.48 60.95 36.80 75.52 44.58 72.26 61.88 60.27
MiLe Loss 68.62 64.17 61.52 39.00 75.41 44.63 72.90 63.61 61.23
1-shot Cross-Entropy Loss 64.13 63.33 58.92 40.40 75.46 48.72 72.90 63.85 60.96
MiLe Loss 65.26 63.93 60.57 38.80 75.46 49.64 72.58 63.93 61.27
5-shot Cross-Entropy Loss 64.22 63.92 60.90 39.60 75.84 51.18 73.60 64.72 61.75
MiLe Loss 66.85 64.58 64.33 41.00 75.14 52.66 74.02 66.06 63.08
@@ 5.3 Performance of Training with More Tokens
MiLe Loss assesses the learning difficulty of each token through information entropy. Intuitively, the more tokens used in model training, the more powerful the Language Model becomes, and the output word distribution becomes more reasonable. Consequently, the assessment of the learning difficulties of tokens becomes more accurate, and thus the MiLe Loss can probably better lead the LM to tackle the bias. To validate that, with limited computational resources, we continue to pre-train the 6.7B model from 100B tokens to 200B tokens, with both Cross-Entropy Loss and MiLe Loss. Their corresponding evaluation results on all benchmarks are reported in the following Table 8. We can see that, when the number of training tokens for the 6.7B models increases to 200B, the models trained with MiLe Loss yield consistent and substantial performance improvements over those trained with Cross-Entropy Loss. Moreover, compared to training with 100B tokens, training with more tokens even helps MiLe Loss to yield LARGER performance improvements. For instance, in the 5-shot setting, with 100B training tokens, the performance improvements gained by MiLe Loss over Cross-Entropy Loss on the 6.7B models is 1.02%. Then by continuing pre-training with more training tokens, the gained improvements increase to 1.33%. The experimental results above demonstrate well that using more tokens and longer pre-training increases the benefits of MiLe Loss.